# Bar Evolution in Edge-on Galaxies: A Demographic Study of Boxy/Peanut Bulges

### v1

@@ Abstract
Methods. We selected a sample of 6684 edge-on galaxies from SDSS DR8 using Galaxy Zoo classifications, requiring a high edge-on probability () and a minimum of 10 independent votes. Two-dimensional image decomposition is performed using GALFIT to obtain structural parameters. Residual images are visually inspected to classify BP/X features into four categories: strong both-sided, both-sided, one-sided, and control (no BP/X). We also estimated stellar mass, distance, and physical size for each galaxy.
Results. Out of 6653 classified galaxies, we identified 1675 (25%) with both-sided BP/X features—545 (8%) strong and 1130 (17%) faint—as well as 1108 (17%) one-sided structures, making up a total of 2783 BP/X-hosting galaxies (42%). One-sided structures, likely signatures of ongoing buckling, are more frequent than strong both-sided bulges across all stellar masses. The fraction of BP/X bulges increases with stellar surface mass density, indicating a connection with bar formation in dense disks. We also find that galaxies with strong BP/X bulges contribute to increased scatter in the stellar mass–size and stellar mass–surface density relations, particularly at higher masses.
@@
Galaxies: formation – Galaxies: evolution1 Introduction
Spiral galaxies with elongated bar-like structures in their center are commonly known as barred galaxies and comprise nearly 30-70% of all spiral galaxies (Aguerri et al., 2009; Masters et al., 2011; D´ıaz-Garc´ıa et al., 2016). Bars play a major role in the evolution of spiral galaxies (Marchuk et al., 2022). They are thought to funnel a large amount of gas into the nuclear region (Shlosman et al., 1989; Boone et al., 2007). This gas helps in the formation of new stars in the central and bar regions (D´ıaz-Garc´ıa et al., 2020; Fraser-McKelvie et al., 2020). The accretion of gas into the nuclear region also feeds the central supermassive black holes (SMBHs) that help the growth of SMBHs. Sometimes the excess gas accretion onto SMBH ignites the nuclear activity in the galaxy (Alonso et al., 2013; Galloway et al., 2015; Silva-Lima et al., 2022; Kataria & Vivek, 2024).
During the growth phase of bars, nuclear stellar disc (NSD) is formed due to intense star formation in the central sub-kpc region (Friedli & Benz, 1993; Heller & Shlosman, 1994; Friedli & Benz, 1995; Martin & Friedli, 1997; Wozniak, 2007; Kim et al., 2011; Cole et al., 2014; Seo et al., 2019; Baba & Kawata, 2020; Sormani et al., 2022). The bar suppresses star formation through out the remainder of its extent (Martin & Friedli, 1997; Spinoso et al., 2016; Khoperskov et al., 2018; Donohoe-Keyes et al., 2019). Therefore, the NSD has younger stars than the age of the bar (Baba & Kawata, 2020) and the BPX bulge contains stars older than the NSD (Gonzalez et al., 2017). Using their numerical simulations, Baba & Kawata (2020) showed that the oldest stellar population in the NSD can be used for dating bar formation time (also see de Sá-Freitas et al., 2023).
@@
There have been a few attempts to study the set of galaxies having BP/X bulges (Ciambur & Graham, 2016; Laurikainen & Salo, 2016; Savchenko et al., 2017). However, it is not easy to detect the presence of a BP/X bulge, as it requires the bulge to be oriented in a perpendicular direction with respect to our line of sight. There has been an attempt to create a catalogue of edge-on galaxies (Kautsch et al., 2006). Additionally, there have been attempts to study the relationship between the development of BP/X bulges and the dark matter halo using numerical simulations (Kumar et al., 2022). Several previous studies (discussed in Section 7) have shown a strong correlation between stellar mass and the presence of BP/X bulges. These studies have been conducted with relatively smaller sample sizes than ours. They concluded that high-mass galaxies are more likely to show the presence of a BP/X bulge than lower-mass galaxies. The gas mass fraction could play a role in determining the way in which the bar develops—either via buckling or symmetric growth—as suggested by Berentzen et al. (2007). However, the current gas mass fraction does not appear to influence whether a BP/X bulge will form (Erwin & Debattista, 2017).
Here, we aim to separate the edge-on galaxies having BP/X bulge from a large sample of edge-on galaxies obtained from the Sloan Digital Sky Survey111SDSS. We classify galaxies having BP/X bulges and study different parameters like masses and sizes of galaxies with or without BP/X bulges. Then we compare their properties to find any speciality in the sample with BP/X bulges. We verified some of the past similar analysis performed previously on relatively smaller sample size. Further, we investigate the effect of BP/X hosting galaxies on the scaling laws, mass-size relation.
@@ 2.1 The data base
The main motivation for choosing minimum 10 votes for each galaxy is to avoid an uncertain classification due to small number of voting. A fraction greater than 0.9 signifies that the most volunteers have identified the galaxy as edge-on. Then we obtain coordinates (RA, Dec) of the selected galaxies from the ”Galaxy” table of the Sloan Digital Sky Survey data release 8 (SDSS DR8)222SDSS DR8. The ”Galaxy” table contains photometric data of all the objects classified as galaxies from the ”PhotoPrimary” table. Though SDSS DR17 is currently available, we use DR8 because ZooSpec table builds on the DR7 and its link is available with DR8 (Aihara et al., 2011). These conditions return us a set of 9068 edge-on galaxies. The SQL search query for this task is attached below.
SELECT g.ra, g.dec, g.sky_r, g.petroRad_r, g.sky_r, g.expPhi_r FROM Galaxy g JOIN zooSpec z ON z.objid = g.objID WHERE z.nvote > 10 AND z.p_edge > 0.90
@@ 2.2 FITS image files
Next, we use the (RA, Dec) of the egde-on galaxies to download the r-band images in FITS (Flexible Image Transport System) format. Our choice of -band images reflects our interest in the stellar light profile of galaxies because -band is less affected by dust extinction and has higher S/N ratio (York et al., 2000), which makes it easier to distinguish other structures from the oval (Yoshino & Yamauchi, 2015). For this purpose, we utilise the module 333Skyview of the python package Astroquery (Ginsburg et al., 2019). Astroquery provide a set of tools for accessing astronomical data from the webs of various surveys. To obtain the complete image of the galaxy, we use the radius of the field to be 1.5 times the Petrosian radius of the galaxy. The pre-computed Petrosian radii of the galaxies are obtained using SDSS SQL search query along with the (RA, Dec) positions as discussed in Section 2.1. For all the galaxies, we used the default Skyview image dimensions, i.e., 300 pixels 300 pixels. Due to our default image dimensions, the pixel size varies among galaxies depending upon the corresponding petrosian radius. Out of 9068 galaxies, we can obtain images for 6684 galaxies because Skyview does not return images for some specific positions. Through a quick visual inspection we removed incomplete images (where only some part of the galaxy is visible) leaving a sample of 6653 galaxies for further analysis. For all the galaxies in our sample we download the corresponding PSF file from SDSS Science Archive Server.
@@ 2.3 Other relevant parameters
Along with obtaining FITS images of the edge-on galaxies, we also complied some relevant parameters which we will use later in this study. These parameters include object ID (objid), redshift (z), absolute magnitudes. For this purpose we make use of (RA, DEC) positions of selected galaxies in SDSS crossID tool 444SDSS crossID. This tool allows users to upload their own table which can be joined with different SDSS tables to extract other information using a similar SQL query as described in Section 2.1. We extract the object ID (objid), absolute magnitude in -, -, & -band from Photoz table, and redshift () from the SpecObjAll table. We also obtain exponential fitting parameters like minor to major axis ratio (), effective radius (Re), position angle, and magnitude from the PhotoObjAll table.
@@ 3.1 Median/Gaussian filter
Using the median or Gaussian filter to smooth the image, and then subtracting these smoothed image from original image (and sometimes adding a factor of residual image back into the original image) to sharp the hidden BP/X structure has been used in the literature (Faúndez-Abans & de Oliveira-Abans, 1998). Therefore, we first started with using the median filter available in the Scipy Python package555Scipy (Virtanen et al., 2020a) to identify the BP/X hosting galaxies. The median filter smooth any image according to input box (or window) size. We give galaxy image and the box size as input parameters. The value of each pixel in the output image is calculated as the median value of the pixels inside the box of the given size around that pixel. Then we subtract the median filtered image from the original image to get the residual image. The residual image shows the salient features hidden near the central part of the galaxy as the outer regions mostly remain the same in both images and get cancel out in the residual image.
@@ 3.2 Ellipse fitting
The subtraction of smooth intensity distribution of galaxy from its observed intensity distribution can also sharpen the hidden BP/X structure. Therefore, we construct smooth intensity distribution of galaxy by fitting iso-intensity ellipses using a Python package Photutils 666ellipse.photutils (Bradley & Developers (2023)). It requires an initial guess for an ellipse geometry to start the fitting. The initial ellipse is described by the user guess for center, semi-major axis, ellipticity, and position angle. We guess initial parameters to perform ellipse fitting, and use the fitted ellipses to model the smooth intensity profile of the galaxy. Then we subtract this model image from observed image to find the residual image. This residual images show the hidden structures behind the smooth intensity distribution of galaxy.
@@
For each galaxy, all these initial guess parameters are compiled into a text file in galfit input format. It also includes galaxy name, output file name, convolution box size among other parameters. Our wrapping algorithm for automation takes care of the variable parameters, e.g., initial guesses, galaxy name, output file name, etc. This fitting analysis provides us with edge-disk parameters which are not available in the original SDSS catalogue. In Fig. 1, we show comparison of exponential disk parameters obtained from SDSS and edge-disk parameters from our galfit decomposition. The left panel compares position angles, and the middle panel compares scale radii. Our analysis shows a good agreement between SDSS and galfit position angles and scale radii as evident from the equality line displayed in left most and middle panels. The SDSS exponential fitting semi-minor to semi-major axis ratio () and galfit edge-disk fitting scale height to scale radius ratio () exhibits linear relation but non-unity slope as shown in the right most panel. We fit the straight line and obtain following relation,
@@ 4 BP/X bulge Classification
After edge-disk fitting using galfit, we visually inspected residual images ( original image model image) of all 6653 galaxies to build a catalogue of galaxies with and without BP/X bulges for detailed study. The sample without BP/X bulges is considered as control sample for comparison with the sample having BP/X bulges. We identify BP/X bulges in different stages/contrast. Following is our classification scheme for BP/X bulges and control sample.
Both-sided:- This class of galaxies show the presence of x-shape or peanut-shape structure at their center in the residual images. The structure can be clearly visible as in Fig. 2(a,b,c) or can be faintly visible as in Fig. 2(d,e,f). We identify 1675 galaxies in this class, which comprise nearly of the total sample.
Strong both-sided:- From the class of both-sided BP/X bulge galaxies, we separate galaxies showing a clear x-shape structures and classify them into strong both-sided sample. These galaxies show a very prominent x-shape structure as in Fig. 2(a,b,c). There are 545 galaxies (nearly ) in this class.
One-sided:- The residual images of these galaxies shows excess intensity on one side of the galactic mid-plane. It seems like the bulge is present on only one side of the galaxy. Fig. 2(g,h,i) is an example of the galaxy in this sample. One of the possible reason for this appearance is the dust obscuration on one side. Additionally, it is also possible that this class potentially contains galaxies in the initial buckling phase. There are 1108 galaxies in this class, which is nearly of the total sample.
Control:- These galaxies show no sign of BP/X bulge formation. Therefore, the model completely fits the original image, leaving no feature in the residual. There are a total of 3870 galaxies in this class. Some of the galaxies in the control sample show presence of other significantly bright structure in the projection, e.g., in the bottom left side of Fig. 2(j,k,l). There are 146 galaxies which show nearby bright structures (likely neighbor galaxy). The remaining 3724 galaxies do not show any neighboring bright objects in the projection. So, nearly of our sample belongs to this class. Note that some of the galaxies showing showing spheroidal structures (classical bulge or end-on projected bars) are also included in the control sample.
@@ 5 Derived Quantities for Galaxies
Along with the parameters obtained from SDSS catalogue and measured using 2D image decomposition using galfit, we derive stellar masses, distances and physical sizes of the galaxies from existing information. The following subsections describes our methodology for measuring these derived properties of galaxies.
@@ 5.1 Stellar masses
After calculating mass-to-light ratio using above described method, we use absolute magnitudes (Mr) to measure the luminosities of galaxies in solar unit by following expression.
where, M⊙ is solar magnitude in . For our analysis, we obtained solar magnitude from Blanton et al. (2003). Next, we measure stellar mass of the galaxies in solar mass unit by multiplying mass-to-light ratio with their luminosities as follows,
@@ 5.2 Angular distances and physical sizes
From the galfit photometric decomposition, we obtain the disk scale height and disk scale radius in pixels. Since we have downloaded the FITS images of galaxies within 3 Petrosian radius aperture with 300 pixels 300 pixels size, the pixel size is simply Petrosian radius / 100 in the same unit as Petrosian radius (here, ). Using the pixel size we calculate the angular size of galaxies in radian. Then we obtained the physical sizes of galaxies by multiplying angular sizes and angular distances.

### v2

@@ Abstract
Methods. We selected a sample of 6684 edge-on galaxies from SDSS DR8 using Galaxy Zoo classifications, requiring a high edge-on probability () and a minimum of 10 independent votes. Two-dimensional image decomposition is performed using GALFIT to obtain structural parameters. Residual images are visually inspected to classify BP/X features into four categories: strong both-sided, both-sided, one-sided, and control (no BP/X). We also estimate stellar mass, distance, and physical size for each galaxy.
Results. Out of 6653 classified galaxies, we identified 1673 (25%) with both-sided BP/X features—504 () strong and 1169 () weak—as well as 1112 () one-sided structures, making up a total of 2785 BP/X-hosting galaxies (). One-sided structures, likely signatures of ongoing buckling, are more frequent than strong both-sided bulges across all stellar masses. The fraction of BP/X bulges increases with stellar surface mass density, indicating a connection with bar formation in dense disks. We also find that galaxies with strong BP/X bulges contribute to increased scatter in the stellar mass–size and stellar mass–surface density relations, particularly at higher masses.
@@
Galaxies: formation – Galaxies: evolution – Galaxies: bulges – Galaxies: structure – Galaxies: photometry – Galaxies: statistics1 Introduction
Spiral galaxies with elongated bar-like structures in their center are commonly known as barred galaxies and comprise nearly 30-70% of all spiral galaxies (Aguerri et al., 2009; Masters et al., 2011; Díaz-García et al., 2016). Bars play a major role in the evolution of spiral galaxies (Marchuk et al., 2022). They are thought to funnel a large amount of gas into the nuclear region (Shlosman et al., 1989; Boone et al., 2007). This gas helps in the formation of new stars in the central and bar regions (Díaz-García et al., 2020; Fraser-McKelvie et al., 2020). The accretion of gas into the nuclear region also feeds the central supermassive black holes (SMBHs) that help the growth of SMBHs. Sometimes the excess gas accretion onto SMBH ignites the nuclear activity in the galaxy (Alonso et al., 2013; Galloway et al., 2015; Silva-Lima et al., 2022; Garland et al., 2024; Kataria & Vivek, 2024).
During the growth phase of bars, nuclear stellar disc (NSD) is formed due to intense star formation in the central sub-kpc region (Friedli & Benz, 1993; Heller & Shlosman, 1994; Friedli & Benz, 1995; Martin & Friedli, 1997; Wozniak, 2007; Kim et al., 2011; Cole et al., 2014; Seo et al., 2019; Baba & Kawata, 2020; Sormani et al., 2022). The bar suppresses star formation through out the remainder of its extent (Martin & Friedli, 1997; Spinoso et al., 2016; Khoperskov et al., 2018; Donohoe-Keyes et al., 2019). Therefore, the NSD has younger stars than the age of the bar (Baba & Kawata, 2020). Using their numerical simulations, Baba & Kawata (2020) showed that the oldest stellar population in the NSD can be used for dating bar formation time (also see de Sá-Freitas et al., 2023).
@@
There have been a few attempts to study the set of galaxies having BP/X bulges (Ciambur & Graham, 2016; Laurikainen & Salo, 2016; Savchenko et al., 2017). However, it is not easy to detect the presence of a BP/X bulge, as it requires the bulge to be oriented in a perpendicular direction with respect to our line of sight. There has been an attempt to create a catalog of edge-on galaxies (Kautsch et al., 2006). Additionally, there have been attempts to study the relationship between the development of BP/X bulges and the dark matter halo using numerical simulations (Kumar et al., 2022). Several previous studies (discussed in Section 7) have shown a strong correlation between stellar mass and the presence of BP/X bulges. These studies have been conducted with relatively smaller sample sizes than ours. They concluded that high-mass galaxies are more likely to show the presence of a BP/X bulge than lower-mass galaxies. The gas mass fraction could play a role in determining the way in which the bar develops—either via buckling or symmetric growth—as suggested by Berentzen et al. (2007). However, the current gas mass fraction does not appear to influence whether a BP/X bulge will form (Erwin & Debattista, 2017).
Here, we aim to separate the edge-on galaxies having BP/X bulge from a large sample of edge-on galaxies obtained from the Sloan Digital Sky Survey. We classify galaxies having BP/X bulges and study different parameters like masses and sizes of galaxies with or without BP/X bulges. Then we compare their properties to find any specialty in the sample with BP/X bulges. We verified some of the past similar analysis performed previously on relatively smaller sample size. Further, we investigate the effect of BP/X hosting galaxies on the scaling laws, mass-size relation.
@@ 2.1 The data base
The main motivation for choosing minimum 10 votes for each galaxy is to avoid an uncertain classification due to small number of voting. A fraction greater than 0.9 signifies that the most volunteers have identified the galaxy as edge-on. Then we obtain coordinates (RA, Dec) of the selected galaxies from the ”Galaxy” table of the Sloan Digital Sky Survey data release 8 (SDSS DR8)111SDSS DR8. The ”Galaxy” table contains photometric data of all the objects classified as galaxies from the ”PhotoPrimary” table. Though SDSS DR17 is currently available, we use DR8 because ZooSpec table builds on the DR7 and its link is available with DR8 (Aihara et al., 2011). These conditions return us a set of 9068 edge-on galaxies. The SQL search query for this task is attached below.
SELECT
g.ra, g.dec, g.sky_r, g.petroRad_r,
g.sky_r, g.expPhi_r
FROM
Galaxy g
JOIN zooSpec z ON z.objid = g.objID
WHERE
z.nvote > 10 AND
z.p_edge > 0.90
@@ 2.2 FITS image files
Next, we use the (RA, Dec) of the egde-on galaxies to download the r-band images in FITS (Flexible Image Transport System) format. Our choice of -band images reflects our interest in the stellar light profile of galaxies because -band is less affected by dust extinction and has higher S/N ratio (York et al., 2000), which makes it easier to distinguish other structures from the oval (Yoshino & Yamauchi, 2015). For this purpose, we utilize the SDSS module 222astroquery.SDSS of the Astroquery (Ginsburg et al., 2019) Python package. Astroquery provide a set of tools for accessing astronomical data from the webs of various surveys. To obtain the complete image of the galaxy, we use the radius of the field to be 1.5 times the Petrosian radius of the galaxy. The pre-computed Petrosian radii of the galaxies are obtained using SDSS SQL search query along with the (RA, Dec) positions as discussed in Section 2.1. For all the galaxies, we used the default SDSS pixel scale, i.e. arcsec/pixel. Out of 9068 galaxies, we can obtain images for 6684 galaxies because SDSS does not return images for some specific positions. Through a quick visual inspection we removed incomplete images (where only some part of the galaxy is visible) leaving a sample of 6653 galaxies for further analysis. For all the galaxies in our sample we download the corresponding PSF file from SDSS Science Archive Server.
@@ 2.3 Other relevant parameters
Along with obtaining FITS images of the edge-on galaxies, we also complied some relevant parameters which we will use later in this study. These parameters include object ID (objid), redshift (z), absolute magnitudes. For this purpose we make use of (RA, DEC) positions of selected galaxies in SDSS crossID tool 333SDSS crossID. This tool allows users to upload their own table which can be joined with different SDSS tables to extract other information using a similar SQL query as described in Section 2.1. We extract the object ID (objid), absolute magnitude in -, -, & -band from Photoz table, and redshift () from the SpecObjAll table. We also obtain exponential fitting parameters like minor to major axis ratio (), effective radius (Re), position angle, and magnitude from the PhotoObjAll table.
@@ 3.1 Median/Gaussian filter
Using the median or Gaussian filter to smooth the image, and then subtracting these smoothed image from original image (and sometimes adding a factor of residual image back into the original image) to sharp the hidden BP/X structure has been used in the literature (Faúndez-Abans & de Oliveira-Abans, 1998). Therefore, we first started with using the median filter available in the Scipy Python package444Scipy (Virtanen et al., 2020a) to identify the BP/X hosting galaxies. The median filter smooth any image according to input box (or window) size. We give galaxy image and the box size as input parameters. The value of each pixel in the output image is calculated as the median value of the pixels inside the box of the given size around that pixel. Then we subtract the median filtered image from the original image to get the residual image. The residual image shows the salient features hidden near the central part of the galaxy as the outer regions mostly remain the same in both images and get cancel out in the residual image.
@@ 3.2 Ellipse fitting
The subtraction of smooth intensity distribution of galaxy from its observed intensity distribution can also sharpen the hidden BP/X structure. Therefore, we construct smooth intensity distribution of galaxy by fitting iso-intensity ellipses using a Python package Photutils 555ellipse.photutils (Bradley & Developers (2023)). It requires an initial guess for an ellipse geometry to start the fitting. The initial ellipse is described by the user guess for center, semi-major axis, ellipticity, and position angle. We guess initial parameters to perform ellipse fitting, and use the fitted ellipses to model the smooth intensity profile of the galaxy. Then we subtract this model image from observed image to find the residual image. This residual images show the hidden structures behind the smooth intensity distribution of galaxy.
@@
For each galaxy, all these initial guess parameters are compiled into a text file in galfit input format. It also includes galaxy name, output file name, convolution box size among other parameters. Our wrapping algorithm for automation takes care of the variable parameters, e.g., initial guesses, galaxy name, output file name, etc. This fitting analysis provides us with edge-disk parameters which are not available in the original SDSS catalog. In Fig. 1, we show comparison of exponential disk parameters obtained from SDSS and edge-disk parameters from our galfit decomposition. The left panel compares position angles, and the middle panel compares scale radii. Our analysis shows a good agreement between SDSS and galfit position angles and scale radii as evident from the equality line displayed in left most and middle panels. The SDSS exponential fitting semi-minor to semi-major axis ratio () and galfit edge-disk fitting scale height to scale radius ratio () exhibits linear relation but non-unity slope as shown in the right most panel. We fit the straight line and obtain following relation,
@@ 4 BP/X bulge classification
After edge-disk fitting using galfit, we visually inspected residual images ( original image model image) of all 6653 galaxies to build a catalog of galaxies with and without BP/X bulges for detailed study. The sample without BP/X bulges is considered as control sample for comparison with the sample having BP/X bulges. We identify BP/X bulges in different stages/contrast. Following is our classification scheme for BP/X bulges and control sample.
Both-sided:- This class of galaxies show the presence of boxy, peanut, or x-shape structure at their center in the residual images. The structure can be clearly visible as in Fig. 2(a,b,c) or can be faintly visible as in Fig. 2(d,e,f). We identify 1673 galaxies in this class, which comprise nearly of the total sample. On the basis of visibility, this class can be further divided into strong and weak both-sided structures. In our analysis, we refer to both-sided as the collection of both strong and weak both-sided structures, unless specified differently.
Strong both-sided:- From the class of both-sided BP/X bulge galaxies, we separate galaxies showing a clear x-shape structures and classify them into strong both-sided sample. These galaxies show a very prominent x-shape structure as in Fig. 2(a,b,c). There are 504 galaxies (nearly ) in this class.
One-sided:- The residual images of these galaxies shows excess intensity on one side of the galactic mid-plane. It seems like the bulge is present on only one side of the galaxy. Fig. 2(g,h,i) is an example of the galaxy in this sample. One of the possible reason for this appearance is the dust obscuration on one side. Additionally, it is also possible that this class potentially contains galaxies in the initial buckling phase. There are 1112 galaxies in this class, which is nearly of the total sample.
Control:- These galaxies show no sign of BP/X bulge formation. Therefore, the model completely fits the original image, leaving no feature in the residual. There are a total of 3868 galaxies in this class. Some of the galaxies in the control sample show presence of other significantly bright structure in the projection, e.g., in the bottom left side of Fig. 2(j,k,l). There are 146 galaxies which show nearby bright structures (likely neighbor galaxy). The remaining 3722 galaxies do not show any neighboring bright objects in the projection. So, nearly of our sample belongs to this class. Note that some of the galaxies showing showing spheroidal structures (classical bulge or end-on projected bars) are also included in the control sample.
For more examples of different categories of BP/X bulges discussed here, we refer the reader to Fig. 8, 9, and 10 in Appendix A. We note that the weak both-sided and one-sided structures may contain some oblate classical bulge structures, which are difficult to distinguish due to spatial resolution limit of the SDSS. We caution the user to consider this caveat when interpreting results for these two classes of BP/X structures in this study.
@@ 5 Derived quantities for galaxies
Along with the parameters obtained from SDSS catalog and measured using 2D image decomposition using galfit, we derive stellar masses, distances and physical sizes of the galaxies from existing information. The following subsections describe our methodology for measuring these derived properties of galaxies.
@@ 5.1 Stellar masses
After calculating mass-to-light ratio using above described method, we use absolute -band magnitudes (Mr) to measure the luminosities of galaxies in solar unit by following expression.
where, M⊙ is solar magnitude in -band. For our analysis, we obtained solar magnitude from Blanton et al. (2003). Next, we measure stellar mass of the galaxies in solar mass unit by multiplying mass-to-light ratio with their luminosities as follows,
@@ 5.2 Angular distances and physical sizes
From the galfit photometric decomposition, we obtain the disk scale height and disk scale radius in pixels. Using the default SDSS pixel size of 0.396 arcsec/pixel, we calculate the angular size of galaxies in radian. Then we obtained the physical sizes of galaxies by multiplying angular sizes and angular distances.